$SPAD/src/input calcprob.input

Timothy Daly December 3, 2016

Abstract

Cover a range of calculus problems

Contents

2

— * —

)set break resume
)spool calcprob.output
)set message test on
)set message auto off
)clear all
--S 1 of 12
solve(3*x-(x-7)=4*x-5,x)
--R
--R   (1)  [x= 6]
--R                          Type: List(Equation(Fraction(Polynomial(Integer))))
--E 1
--S 2 of 12
solve(4*x-3*y=9,y)::List Equation Polynomial Fraction Integer
--R
--R            4
--R   (2)  [y= - x - 3]
--R            3
--R                          Type: List(Equation(Polynomial(Fraction(Integer))))
--E 2
--S 3 of 12
solve(A*x+B*y=C,y)
--R
--R            - A x + C
--R   (3)  [y= ---------]
--R                B
--R                          Type: List(Equation(Fraction(Polynomial(Integer))))
--E 3
--S 4 of 12
m:=3*x-4*(x-(2/3)*y)=(4/5)*x-(7*y+3)
--R
--R        8               4
--R   (4)  - y - x= - 7y + - x - 3
--R        3               5
--R                                Type: Equation(Polynomial(Fraction(Integer)))
--E 4
--S 5 of 12
n:=solve(m*15,y)
--R
--R            27x - 45
--R   (5)  [y= --------]
--R               145
--R                          Type: List(Equation(Fraction(Polynomial(Integer))))

--E 5
--S 6 of 12
p:=n.1*145-27*x
--R
--R   (6)  145y - 27x= - 45
--R                                Type: Equation(Fraction(Polynomial(Integer)))
--E 6
--S 7 of 12
(x1,y1):=(-3,-8)
--R
--R   (7)  - 8
--R                                                                Type: Integer
--E 7
--S 8 of 12
(x2,y2):=(-6,2)
--R
--R   (8)  2
--R                                                        Type: PositiveInteger
--E 8
--S 9 of 12
m:=(y2-y1)/(x2-x1)
--R
--R          10
--R   (9)  - --
--R           3
--R                                                      Type: Fraction(Integer)
--E 9
--S 10 of 12
solve(y1=m*x1+b,b)
--R
--R   (10)  [b= - 18]
--R                          Type: List(Equation(Fraction(Polynomial(Integer))))
--E 10
--S 11 of 12
b:=-18
--R
--R   (11)  - 18
--R                                                                Type: Integer
--E 11
--S 12 of 12
y=m*x+b
--R
--R              10

--R   (12)  y= - -- x - 18
--R               3
--R                                Type: Equation(Polynomial(Fraction(Integer)))
--E 12
)spool
)lisp (bye)

———-

References^{[1] nothing}

6